OpenStudy (republican31):

I need a formula to solve for s. r = 2s - 8.

OpenStudy (anonymous):

To get s, you should add 8 to both sides r + 8 = 2s NOw if you divide by 2, you get s by itself r/2 + 4 = s

OpenStudy (the_fizicx99):

\[R = 2s -8\] \[R + 8 = 2s\] \[\frac{ R + 8}{ 2 } = S\]